# Joint Modeling of Multivariate Longitudinal and Survival Outcomes with the R package INLAjoint

### v1

@@ Abstract
This paper introduces the R package INLAjoint, designed as a toolbox for fitting a diverse range of regression models addressing both longitudinal and survival outcomes. INLAjoint relies on the computational efficiency of the integrated nested Laplace approximations methodology, an efficient alternative to Markov chain Monte Carlo for Bayesian inference, ensuring both speed and accuracy in parameter estimation and uncertainty quantification. The package facilitates the construction of complex joint models by treating individual regression models as building blocks, which can be assembled to address specific research questions. Joint models are relevant in biomedical studies where the collection of longitudinal markers alongside censored survival times is common. They have gained significant interest in recent literature, demonstrating the ability to rectify biases present in separate modeling approaches such as informative censoring by a survival event or confusion bias due to population heterogeneity. We provide a comprehensive overview of the joint modeling framework embedded in INLAjoint with illustrative examples. Through these examples, we demonstrate the practical utility of INLAjoint in handling complex data scenarios encountered in biomedical research.
@@
$*{}^{*}$orrespondence: Denis Rustand, KAUST, Email: denis@rustand.fr
@@ 3.5 S3 methods
While it is possible to use the predict function to predict the longitudinal and survival models trajectories over new individuals or forecast for some known individuals, it also allows to predict the average trajectory for any outcome conditional on covariates. In this case, the new data should only consist of one line that gives the value of covariates and the longitudinal outcomes values should be set to NA. The predict function will automatically assume the average trajectory of each outcome conditional on the given covariates, and will quantify uncertainty through sampling where the marginal distribution of random effects is used (instead of the distribution conditional on observed longitudinal measurements when provided) which allows to display the average trajectories for inference purposes. This is very useful for complex models as it is common to use models that allow for complex trajectories (e.g., fixed and random effects on splines or complex interactions), where the value of parameters is difficult to directly interpret. With this feature, one can compute the average trajectory conditional on treatment and compare the longitudinal and survival profiles of each treatment line visually for example. This feature is illustrated in the second usage example (Section 4).
@@ 4 Usage Examples
We illustrate the usage of INLAjoint through two examples of multivariate joint modeling. The first one shows how to define a joint model with 2 longitudinal outcomes and a survival outcome to introduce basic concepts of INLAjoint’s usage and compare with alternative software. The second example illustrates advanced features in a model that includes 7 longitudinal outcomes and two competing risks of events, this complex joint model cannot be fitted with alternative software in the literature at the moment. All the computations are done over 12 Intel Xeon Gold 6248 2.50GHz CPUs. We used R version 4.3.2, INLAjoint version 24.2.4, INLA version 24.02.09, JMbayes2 version 0.4-5 and rstanarm version 2.32.1.
@@ 4.2 Example 1: Joint model for two longitudinal outcomes and a terminal event
In this first example, we model two longitudinal outcomes with a lognormal mixed effects model and a Poisson mixed effects model, both including fixed effects for the intercept, slope and drug as well as correlated random intercepts. Let $Y_{i1}(t)$ denote the value of SGOT (i.e., aspartate aminotransferase) for individual $i$ at time $t$ and $Y_{i2}(t)$ the corresponding value for the platelet counts. Assuming $k=1,2$ for SGOT and platelet, respectively, the fixed intercept is denoted by $\beta_{k0}$, the fixed slope by $\beta_{k1}$, the effect of drug at baseline by $\beta_{k2}$ and the effect of drug over time by $\beta_{k3}$. Each longitudinal submodel includes a Gaussian random intercept with variance $\sigma^{2}_{k}$ and their covariance is denoted $\sigma_{12}$. The residual error term for the lognormal observations is denoted by $\varepsilon_{i1}(t)$, which is assumed Gaussian with mean 0 and standard deviation $\sigma_{\varepsilon}$. The risk of the composite event (i.e., death or transplantation) for individual $i$ at time $t$ is denoted by $\lambda_{i}(t)$ and defined by a baseline risk $\lambda_{0}(t)$ and a linear predictor including a fixed effect of drug $\gamma_{1}$ and the effect of the shared linear predictor from the two longitudinal submodels $\varphi_{1}$ and $\varphi_{2}$. The purpose of this example is to compare INLAjoint’s fit with the two available alternatives in R able to fit this model, JMbayes2 and rstanarm, in order to highlight INLAjoint’s computational speed and accuracy. We keep the default priors for all packages despite their differences as the intention is to compare packages as they are meant to be used. Simulation studies evaluated the differences in the case of matching priors between INLAjoint and rstanarm (Rustand et al., 2023a, ). Note that JMbayes2 and rstanarm use LME from univariate models (i.e., mixed effects and proportional hazards models) to define initial values and JMbayes2 also uses those univariate models to define data-driven prior distributions for the joint model fit. The three packages assume Gaussian priors for the fixed effects, where rstanarm assume a variance of 2.5 while INLAjoint assume a variance of 100. For the random effects covariance matrix, rstanarm uses the LKJ prior while JMbayes2 and INLAjoint use inverse-Wishart. More details about the priors and initial values used with INLAjoint can be displayed from a fitted model with the function priors.used(MODEL), where MODEL is the name of the object that contains the fitted model. Moreover, the 3 methods use different approximation of the baseline hazard, rstanarm uses cubic B-splines for the log baseline hazard with 2 internal knots while JMbayes2 uses quadratic B-splines with 9 internal knots. With INLAjoint, we illustrate the data augmentation of a Weibull baseline hazard where the follow-up is decomposed into 15 intervals with equidistant nodes (betwen 0 and maximum observed time). Therefore, the comparison will not focus on the baseline hazard estimation as it is handled differently for each package. The default number of MCMC chains and iterations is used with JMbayes2 (i.e., 6500 iterations and 3 chains for this model) but we increased the default number of iterations with rstanarm to 6000 with 4 chains instead of the default 2000 iterations with 4 chains in order to reach proper convergence as results with less iterations were unstable. The fitted model is defined as follows:
@@ 4.3 Example 2: Joint model for 7 longitudinal outcomes and 2 competing risks of event
Longitudinal outcome (L1, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L1         -0.0783 0.0519    -0.1800  -0.0783     0.0233f1year_L1             1.9711 0.1380     1.7006   1.9711     2.2415f2year_L1             2.0434 0.1879     1.6752   2.0434     2.4117Res. err. (variance)  0.0741 0.0030     0.0683   0.0740     0.0802Random effects variance-covariance (L1)                         mean     sd 0.025quant 0.5quant 0.975quantIntercept_L1           0.8119 0.0780     0.6785   0.8075     0.9824f1year_L1              3.9666 0.5683     3.0664   3.9021     5.2616f2year_L1              3.5344 0.7924     2.1974   3.4765     5.2455Intercept_L1:f1year_L1 0.5931 0.1748     0.2756   0.5830     0.9776Intercept_L1:f2year_L1 0.6012 0.2826     0.1166   0.5817     1.2204f1year_L1:f2year_L1    1.9691 0.6056     1.0007   1.9116     3.3348Longitudinal outcome (L2, poisson)                    mean     sd 0.025quant 0.5quant 0.975quantIntercept_L2      5.4488 0.0999     5.2530   5.4488     5.6446year_L2          -0.1166 0.0496    -0.2139  -0.1166    -0.0193drug_L2          -0.0507 0.1309    -0.3072  -0.0507     0.2058sex_L2            0.0815 0.1051    -0.1246   0.0815     0.2875year:drug_L2     -0.0412 0.0651    -0.1688  -0.0412     0.0864year:sex_L2       0.0592 0.0521    -0.0429   0.0592     0.1614drug:sex_L2      -0.0221 0.1389    -0.2944  -0.0221     0.2502year:drug:sex_L2  0.0027 0.0691    -0.1327   0.0027     0.1381Random effects variance-covariance (L2)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L2          0.1491 0.0119     0.1273   0.1486     0.1743year_L2               0.0310 0.0041     0.0241   0.0307     0.0399Intercept_L2:year_L2 -0.0049 0.0042    -0.0138  -0.0047     0.0031Longitudinal outcome (L3, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L3          0.1127 0.0496     0.0155   0.1127     0.2098year_L3              -0.0123 0.0048    -0.0217  -0.0123    -0.0028Res. err. (variance)  0.3036 0.0105     0.2832   0.3035     0.3246Random effects variance-covariance (L3)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L3 0.6602 0.0602     0.5515    0.657     0.7878Longitudinal outcome (L4, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L4          0.2678 0.0463     0.1771   0.2678     0.3585year_L4              -0.1473 0.0060    -0.1591  -0.1473    -0.1355Res. err. (variance)  0.4878 0.0170     0.4551   0.4877     0.5218Random effects variance-covariance (L4)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L4 0.5032 0.0499     0.4145   0.4999     0.6104Longitudinal outcome (L5, binomial)                mean     sd 0.025quant 0.5quant 0.975quantIntercept_L5 -3.7897 0.1849    -4.1521  -3.7897    -3.4273year_L5       0.2256 0.0325     0.1620   0.2256     0.2893Random effects variance-covariance (L5)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L5 3.7995 0.8292     2.4843   3.6897      5.732Longitudinal outcome (L6, binomial)                mean     sd 0.025quant 0.5quant 0.975quantIntercept_L6 -1.4105 0.1648    -1.7334  -1.4105    -1.0876year_L6       0.1302 0.0252     0.0809   0.1302     0.1796Random effects variance-covariance (L6)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L6 5.2765 0.6945     3.9942   5.2543     6.7166Longitudinal outcome (L7, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L7         -0.2216 0.0449    -0.3096  -0.2216    -0.1336year_L7               0.1086 0.0066     0.0956   0.1086     0.1216Res. err. (variance)  0.6042 0.0211     0.5638   0.6038     0.6468Random effects variance-covariance (L7)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L7 0.4186 0.0465     0.3357   0.4157     0.5185Association longitudinal - survival             mean     sd 0.025quant 0.5quant 0.975quantCV_L1_S1   1.3960 0.1224     1.1561   1.3956     1.6381CV_L2_S1  -0.6470 0.1432    -0.9332  -0.6455    -0.3696CV_L2_S2  -0.4818 0.3086    -1.0898  -0.4817     0.1253CV_L3_S2   0.6404 0.3009     0.0466   0.6409     1.2313SRE_L4_S2 -0.6304 0.3611    -1.3499  -0.6275     0.0720SRE_L5_S1  0.1393 0.0655     0.0083   0.1400     0.2662CV_L6_S2   0.0770 0.1016    -0.1201   0.0760     0.2800CS_L6_S2  -0.1218 0.4358    -0.9856  -0.1197     0.7302CS_L7_S1  -0.0449 0.6640    -1.3783  -0.0360     1.2359log marginal-likelihood (integration)    log marginal-likelihood (Gaussian)                            -94888.52                             -94861.59Deviance Information Criterion:  -108645.6Widely applicable Bayesian information criterion:  -107395.6Computation time: 293.12 seconds
@@
Now, we illustrate how to use the predict function. We need to define a new dataset for which we want predictions. Here for simplicity we use data of an individual from the training set but treated as a new individual. In order to illustrate the possibility to deal with dynamic predictions, we define two landmark points, such that the first prediction is based on the first 3 longitudinal measurements while the second prediction is updated to take into account the entire history of longitudinal markers for this individual (i.e., 12 observations). Moreover, we add a line at the end of the new data for predictions where the covariates values are set to 0 (i.e., reference individual) and the outcome values are set to NA. This will compute predictions of the longitudinal and survival components for the reference individual. By adding a similar line where the variable drug is set to 1 for example, we could compare the average trajectories conditional on drug. This feature can be very useful for inference purposes as it allows to compute the average profiles conditional on covariates values, since it becomes difficult to interpret parameters when complex components are involved such as splines or multiple interactions and it is often easier to visually depict the trajectories of longitudinal and survival components. We set the horizon of the predictions to 14 years.
R> NewDat <- pbc2_2[pbc2_2\$id==25,]R> NewDat\$id <- 2R> NewDat <- rbind(NewDat[1:3, ], NewDat, c(3, 0, 0, 0, rep(NA, 7)))R> NewDat\$id[1:3] <- 1R> HRZ <- 14R> print(NewDat, digits=1, row.names=F)
 id drug sex year ascites spiders log_serBilir albumin log_SGOT platelets log_prothrombin  1    0   1  0.0       0       0         -0.9    1.41     0.03       324            0.31  1    0   1  0.5       0       0         -1.2    1.07     2.05       285           -0.54  1    0   1  1.1       0       0         -1.0    0.62    -1.59       264           -0.45  2    0   1  0.0       0       0         -0.9    1.41     0.03       324            0.31  2    0   1  0.5       0       0         -1.2    1.07     2.05       285           -0.54  2    0   1  1.1       0       0         -1.0    0.62    -1.59       264           -0.45  2    0   1  2.1       0       0         -1.2    0.87    -0.31       219           -0.72  2    0   1  3.1       0       0         -0.9    0.44    -0.22       188            1.31  2    0   1  4.1       0       0         -1.4    0.84    -1.46       150           -1.38  2    0   1  5.1       0       0         -1.2    0.34    -0.37       209           -0.90  2    0   1  6.1       0       1         -1.4   -0.46    -0.55       208           -1.09  2    0   1  7.0       0       1         -1.2    0.14    -1.52       170           -1.00  2    0   1  8.1       0       1         -1.0    0.12    -0.68       110           -0.19  2    0   1  9.0       0       0         -1.2   -0.04    -0.81        98           -0.02  2    0   1 10.0       0       1         -0.9   -1.11    -0.60       148           -0.19  3    0   0  0.0      NA      NA           NA      NA       NA        NA              NA
@@
The plot is given in Figure 7. The cumulative incidence function that quantifies the probability of each competing event conditional on survival from all events is by default evaluated from the last longitudinal observation time but it is possible to set a starting time for survival predictions by setting the argument Csurv equal to the desired value. We can see in this plot that when many longitudinal observations are provided (i.e., landmark with 12 observations per longitudinal marker), the uncertainty tends to be reduced compared to predictions with few observations (i.e., landmark with 3 observations per longitudinal markers) where uncertainty is closer to the reference individual without longitudinal information.

### v2

@@ Abstract
{adjustwidth}
40pt40pt This paper introduces the R package INLAjoint, designed as a toolbox for fitting a diverse range of regression models addressing both longitudinal and survival outcomes. INLAjoint relies on the computational efficiency of the integrated nested Laplace approximations methodology, an efficient alternative to Markov chain Monte Carlo for Bayesian inference, ensuring both speed and accuracy in parameter estimation and uncertainty quantification. The package facilitates the construction of complex joint models by treating individual regression models as building blocks, which can be assembled to address specific research questions. Joint models are relevant in biomedical studies where the collection of longitudinal markers alongside censored survival times is common. They have gained significant interest in recent literature, demonstrating the ability to rectify biases present in separate modeling approaches such as informative censoring by a survival event or confusion bias due to population heterogeneity. We provide a comprehensive overview of the joint modeling framework embedded in INLAjoint with illustrative examples. Through these examples, we demonstrate the practical utility of INLAjoint in handling complex data scenarios encountered in biomedical research.
@@
{adjustwidth}20pt120pt Correspondence: Denis Rustand, KAUST, Email: denis@rustand.fr
@@ 3.5 S3 methods
While it is possible to use the predict function to predict the longitudinal and survival models trajectories over new individuals or forecast for some known individuals, it also allows to predict the average trajectory for any outcome conditional on covariates. In this case, the new data should only consist of one line that gives the value of covariates and the longitudinal outcomes values should be set to NA. The predict function will automatically assume the average trajectory of each outcome conditional on the given covariates, and will quantify uncertainty through sampling where the marginal distribution of random effects is used (instead of the distribution conditional on observed longitudinal measurements when provided) which allows to display the average trajectories for inference purposes. This is very useful for complex models as it is common to use models that allow for complex trajectories (e.g., fixed and random effects on splines or complex interactions), where the value of parameters is difficult to directly interpret. With this feature, one can compute the average trajectory conditional on treatment and compare the longitudinal and survival profiles of each treatment line visually for example. This feature is illustrated in the second usage example (Section 4). Note that predictions with INLAjoint are fully Bayesian, meaning that all the parameters are sampled (i.e., including hyperparameters) to assess uncertainty.
@@ 4 Usage Examples
We illustrate the usage of INLAjoint through two examples of multivariate joint modeling. The first one shows how to define a joint model with 2 longitudinal outcomes and a survival outcome to introduce basic concepts of INLAjoint’s usage and compare with alternative software. The second example illustrates advanced features in a model that includes 7 longitudinal outcomes and two competing risks of events, this complex joint model cannot be fitted with alternative software in the literature at the moment. All the computations are done over 12 Intel Xeon Gold 6248 2.50GHz CPUs. We used R version 4.3.2, INLAjoint version 24.3.25, INLA version 24.03.09, JMbayes2 version 0.4-5 and rstanarm version 2.32.1.
@@ 4.2 Example 1: Joint model for two longitudinal outcomes and a terminal event
In this first example, we model two longitudinal outcomes with a lognormal mixed effects model and a Poisson mixed effects model, both including fixed effects for the intercept, slope and drug as well as correlated random intercepts. Let $Y_{i1}(t)$ denote the value of SGOT (i.e., aspartate aminotransferase) for individual $i$ at time $t$ and $Y_{i2}(t)$ the corresponding value for the platelet counts. Assuming $k=1,2$ for SGOT and platelet, respectively, the fixed intercept is denoted by $\beta_{k0}$, the fixed slope by $\beta_{k1}$, the effect of drug at baseline by $\beta_{k2}$ and the effect of drug over time by $\beta_{k3}$. Each longitudinal submodel includes a Gaussian random intercept with variance $\sigma^{2}_{k}$ and their covariance is denoted $\sigma_{12}$. The residual error term for the lognormal observations is denoted by $\varepsilon_{i1}(t)$, which is assumed Gaussian with mean 0 and standard deviation $\sigma_{\varepsilon}$. The risk of the composite event (i.e., death or transplantation) for individual $i$ at time $t$ is denoted by $\lambda_{i}(t)$ and defined by a baseline risk $\lambda_{0}(t)$ and a linear predictor including a fixed effect of drug $\gamma_{1}$ and the effect of the shared linear predictor from the two longitudinal submodels $\varphi_{1}$ and $\varphi_{2}$. The purpose of this example is to compare INLAjoint’s fit with the two available alternatives in R able to fit this model, JMbayes2 and rstanarm, in order to highlight INLAjoint’s computational speed and accuracy. We keep the default priors for all packages despite their differences as the intention is to compare packages as they are meant to be used. Simulation studies evaluated the differences in the case of matching priors between INLAjoint and rstanarm (Rustand et al., 2023a, ). Note that JMbayes2 and rstanarm use LME from univariate models (i.e., mixed effects and proportional hazards models) to define initial values and JMbayes2 also uses those univariate models to define data-driven prior distributions for the joint model fit. The three packages assume Gaussian priors for the fixed effects, where rstanarm assume a variance of 2.5 while INLAjoint assume a variance of 100. For the random effects covariance matrix, rstanarm uses the LKJ prior while JMbayes2 and INLAjoint use inverse-Wishart. More details about the priors and initial values used with INLAjoint can be displayed from a fitted model with the function inla.priors.used(MODEL), where MODEL is the name of the object that contains the fitted model. Moreover, the 3 methods use different approximation of the baseline hazard, rstanarm uses cubic B-splines for the log baseline hazard with 2 internal knots while JMbayes2 uses quadratic B-splines with 9 internal knots. With INLAjoint, we illustrate the data augmentation of a Weibull baseline hazard where the follow-up is decomposed into 15 intervals with equidistant nodes (betwen 0 and maximum observed time). Therefore, the comparison will not focus on the baseline hazard estimation as it is handled differently for each package. The default number of MCMC chains and iterations is used with JMbayes2 (i.e., 6500 iterations and 3 chains for this model) but we increased the default number of iterations with rstanarm to 6000 with 4 chains instead of the default 2000 iterations with 4 chains in order to reach proper convergence as results with less iterations were unstable. The fitted model is defined as follows:
@@ 4.3 Example 2: Joint model for 7 longitudinal outcomes and 2 competing risks of event
Longitudinal outcome (L1, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L1         -0.0784 0.0518    -0.1800  -0.0784     0.0232f1year_L1             1.9748 0.1386     1.7030   1.9748     2.2465f2year_L1             2.0532 0.1865     1.6877   2.0532     2.4186Res. err. (variance)  0.0741 0.0030     0.0684   0.0740     0.0802Random effects variance-covariance (L1)                         mean     sd 0.025quant 0.5quant 0.975quantIntercept_L1           0.8161 0.0784     0.6803   0.8102     0.9845f1year_L1              3.9244 0.5821     2.9605   3.8636     5.2134f2year_L1              3.5459 0.8087     2.1951   3.4625     5.3296Intercept_L1:f1year_L1 0.5876 0.1700     0.2851   0.5747     0.9398Intercept_L1:f2year_L1 0.5966 0.2836     0.0649   0.5777     1.1698f1year_L1:f2year_L1    1.9361 0.6285     0.9121   1.8699     3.3822Longitudinal outcome (L2, poisson)                    mean     sd 0.025quant 0.5quant 0.975quantIntercept_L2      5.4487 0.1000     5.2527   5.4487     5.6448year_L2          -0.1166 0.0497    -0.2140  -0.1166    -0.0191drug_L2          -0.0506 0.1310    -0.3074  -0.0506     0.2062sex_L2            0.0814 0.1053    -0.1249   0.0814     0.2878year:drug_L2     -0.0410 0.0652    -0.1687  -0.0410     0.0868year:sex_L2       0.0595 0.0522    -0.0429   0.0595     0.1618drug:sex_L2      -0.0222 0.1391    -0.2948  -0.0222     0.2504year:drug:sex_L2  0.0028 0.0691    -0.1328   0.0028     0.1383Random effects variance-covariance (L2)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L2          0.1492 0.0118     0.1272   0.1487     0.1741year_L2               0.0309 0.0042     0.0236   0.0306     0.0394Intercept_L2:year_L2 -0.0041 0.0050    -0.0142  -0.0041     0.0061Longitudinal outcome (L3, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L3          0.1127 0.0495     0.0156   0.1127     0.2098year_L3              -0.0123 0.0048    -0.0217  -0.0123    -0.0029Res. err. (variance)  0.3034 0.0107     0.2829   0.3032     0.3250Random effects variance-covariance (L3)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L3 0.6598 0.0655     0.5417   0.6561      0.799Longitudinal outcome (L4, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L4          0.2678 0.0463     0.1770   0.2678     0.3586year_L4              -0.1474 0.0060    -0.1592  -0.1474    -0.1356Res. err. (variance)  0.4878 0.0170     0.4555   0.4874     0.5224Random effects variance-covariance (L4)               mean    sd 0.025quant 0.5quant 0.975quantIntercept_L4 0.5012 0.052     0.4068   0.4985     0.6112Longitudinal outcome (L5, binomial)                mean     sd 0.025quant 0.5quant 0.975quantIntercept_L5 -3.7768 0.1841    -4.1377  -3.7768    -3.4159year_L5       0.2243 0.0324     0.1608   0.2243     0.2878Random effects variance-covariance (L5)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L5 3.4529 0.6392     2.2997    3.426     4.7976Longitudinal outcome (L6, binomial)              mean     sd 0.025quant 0.5quant 0.975quantIntercept_L6 -1.41 0.1647    -1.7328    -1.41    -1.0872year_L6       0.13 0.0252     0.0807     0.13     0.1794Random effects variance-covariance (L6)              mean     sd 0.025quant 0.5quant 0.975quantIntercept_L6 5.269 0.9292     3.6123   5.2195      7.251Longitudinal outcome (L7, gaussian)                        mean     sd 0.025quant 0.5quant 0.975quantIntercept_L7         -0.2215 0.0451    -0.3098  -0.2215    -0.1332year_L7               0.1087 0.0066     0.0957   0.1087     0.1217Res. err. (variance)  0.6034 0.0213     0.5630   0.6030     0.6465Random effects variance-covariance (L7)               mean     sd 0.025quant 0.5quant 0.975quantIntercept_L7 0.4202 0.0479     0.3333   0.4178     0.5215Survival outcome (S1)                              mean     sd 0.025quant 0.5quant 0.975quantBaseline risk (variance)_S1 0.0299 0.0421     0.0013    0.016     0.1462Survival outcome (S2)                            mean     sd 0.025quant 0.5quant 0.975quantBaseline risk (variance)_S2 0.17 0.2013     0.0125   0.1055     0.7307Association longitudinal - survival             mean     sd 0.025quant 0.5quant 0.975quantCV_L1_S1   1.4001 0.1201     1.1597   1.4014     1.6328CV_L2_S1  -0.6632 0.1954    -1.0466  -0.6637    -0.2772CV_L2_S2  -0.5142 0.2752    -1.0439  -0.5183     0.0396CV_L3_S2   0.6299 0.4178    -0.1868   0.6279     1.4584SRE_L4_S2 -0.6602 0.3501    -1.3400  -0.6634     0.0385SRE_L5_S1  0.1403 0.0719    -0.0038   0.1412     0.2792CV_L6_S2   0.0564 0.1091    -0.1634   0.0581     0.2663CS_L6_S2  -0.0125 1.1612    -2.3159  -0.0066     2.2564CS_L7_S1   0.1417 1.1917    -2.1448   0.1218     2.5469log marginal-likelihood (integration)    log marginal-likelihood (Gaussian)                            -94885.27                             -94858.34Deviance Information Criterion:  -108646.8Widely applicable Bayesian information criterion:  -89290.45Computation time: 116.87 seconds
@@
Now, we illustrate how to use the predict function. We need to define a new dataset for which we want predictions. Here for simplicity we use data of an individual from the training set but treated as a new individual. In order to illustrate the possibility to deal with dynamic predictions, we define two landmark points, such that the first prediction is based on the first 2 longitudinal measurements while the second prediction is updated to take into account the entire history of longitudinal markers for this individual (i.e., 6 observations). Moreover, we add a line at the end of the new data for predictions where the covariates values are set to 0 (i.e., reference individual) and the outcome values are set to NA. This will compute predictions of the longitudinal and survival components for the reference individual. By adding a similar line where the variable drug is set to 1 for example, we could compare the average trajectories conditional on drug. This feature can be very useful for inference purposes as it allows to compute the average profiles conditional on covariates values, since it becomes difficult to interpret parameters when complex components are involved such as splines or multiple interactions and it is often easier to visually depict the trajectories of longitudinal and survival components. We set the horizon of the predictions to 14 years.
R> NewDat <- pbc2_2[pbc2_2\$id==139,]R> NewDat\$id <- 2R> NewDat <- rbind(NewDat[1:2, ], NewDat, c(3, 0, 0, 0, rep(NA, 7)))R> NewDat\$id[1:2] <- 1R> HRZ <- 14R> print(NewDat, digits=1, row.names=F)
 id drug sex year ascites spiders log_serBilir albumin log_SGOT platelets log_prothrombin  1    0   1  0.0       0       0        -0.46   -0.16    -0.19       142             0.6  1    0   1  0.5       0       0        -0.74   -0.36     0.08       120             0.6  2    0   1  0.0       0       0        -0.46   -0.16    -0.19       142             0.6  2    0   1  0.5       0       0        -0.74   -0.36     0.08       120             0.6  2    0   1  1.0       0       0        -0.74    0.02    -0.36       121            -0.1  2    0   1  2.0       1       0        -0.31   -0.44    -0.51        81             0.6  2    0   1  2.9       0       0         0.03   -0.08    -0.42        64             1.5  2    0   1  5.0       0       0         0.81   -1.25    -0.22        59             2.3  3    0   0  0.0      NA      NA           NA      NA       NA        NA              NA
@@
The plot is given in Figure 7. The cumulative incidence function that quantifies the probability of each competing event conditional on survival from all events is by default evaluated from the last longitudinal observation time but it is possible to set a starting time for survival predictions by setting the argument Csurv equal to the desired value. We can see in this plot that when many longitudinal observations are provided (i.e., landmark with 6 observations per longitudinal marker), the uncertainty tends to be reduced compared to predictions with few observations (i.e., landmark with 2 observations per longitudinal markers) where uncertainty is closer to the reference individual without longitudinal information.